Fermi National Laboratory


Accelerator Update

The integrated luminosity for last week was 6.8 pb-1

Wednesday August 13
The day shift started with store #2898 colliding, stacking, and beam to Miniboone.

Switchyard, Slipstacking, and 2.5 MHZ studies occurred throughout the shift.

At 11:32 AM a quench at C36L killed the store. The cause of the quench was traced to bad bearings on the C3 cold compressor.

CDF and D0 took advantage of this downtime to make controlled access into their experimental halls.

At 2:40 PM the C3 compressor bearing change was complete and cool down was beginning.

Switchyard, Slipstacking, and 2.5 MHZ studies were all completed by 4:30 PM.

At 5:00 PM D0 made an access to replace a power supply that powers half of the calorimeter.

Tevatron (Tev) checkout was complete by 6:00 PM.

Shot setup began at 10:18 PM.

Thursday August 14
Store #2904 was colliding by 12:34 AM with an initial average luminosity of 41.02e30.

We then returned to stacking.

Store #2904 was lost at 9:12 AM when a Tev RF (TRF3) station tripped causing a quench at B0.

Problems due to a database crash on the midnight shift delayed Tev startup.

Shot setup started at 2:19 PM.

Store #2908 was colliding by 4:52 PM with an initial average luminosity of 39.27e30.

We then returned to stacking.

The store was lost at 9:44 PM when a quench occurred at BA.

The quench was caused by a bad VFC card at BAD1. The card was changed out before the end of shift.

Friday August 13
Shot setup was underway by 2:35 AM, with store #2910 colliding by 5:00 AM with an initial average luminosity of 38.98e30.

We then returned to stacking.

Frequent trips on Linac RF station #2 may require repair work on the dayshift.

Today's Plan
Stack and shot and continue sending beam to MiniBoone.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/15/2003   email Fermilab